# From Fibers to Cells: Fourier-Based Registration Enables Virtual Cresyl Violet Staining From 3D Polarized Light Imaging

### v1

@@ 3.5 Reliability of predicted cell instances across cell sizes
Figure 10: Reliability of predicted cell instances increases with their size. We report F1 scores between detected cell instances by a CPN model in the predicted and target stainings. F1 scores are reported for intervals of the segmented in-plane cell body area. Each bar represents the average over four independently trained models. Error bars show standard deviation.
@@
Larger cells are expected to be more pronounced in 3D-PLI parameter maps compared to smaller cells that may be overshadowed by other tissue components such as nerve fibers. To quantify the reliability of predicted cells in the virtual staining in relation to their size, we compute F1 scores for multiple bins of 2D in-plane cell sizes. Obtaining F1 scores for each bin requires computation of true and false positives, as well as false negatives through matching of predicted and target cells with a minimum IoU threshold of 0.3. We modify the computation of true and false positives by matching only predicted cells within that range with all cells in the target image. To count false negatives, target cells within that range are matched with all cells in the prediction. Cell sizes are corrected with the 2D shrinkage factor of 0.97 estimated in Sec. 2.1.2, assuming homogeneous shrinkage of cells and surrounding tissue.
Results in Fig. 10 show that smaller cells < 50 µm have much lower F1 scores, below 20.0, with all methods. With increasing 2D cell size, they can be identified more accurately. The Gram+Reg model shows highest F1 scores across all cell sizes and smallest variation between four independently trained models, measured as standard deviation.
@@ 4 Discussion
We introduced an online registration head capable of approximating smooth nonlinear deformations, given sufficiently small patch sizes for training. Its accuracy is inherently influenced by the choice of the registration metric. We identified MSE in the Fourier domain through multiple applications of cross-correlation as superior to other typically used metrics in Fourier-based image registration, including PC. This is a notable observation, as PC is a popular choice due to its robustness to intensity variations and frequency-dependent noise Tong et al., (2019). However, during early training stages, model predictions are often blurred. The network tends to learn coarse, low-frequency structures first. Additionally, random weight initialization introduces noise in the first steps. MSE appears more resilient to such blur and frequency-independent noise, leading to more accurate predictions (Fig. 6). This behavior can be explained by PC’s reliance on strong high-frequency spectral peaks, which are diminished by blur Pedone et al., (2013). MSE, in contrast, maintains sensitivity across the frequency spectrum and can better handle smooth, low-detail inputs that dominate during initial training.
@@
The applied pixel-wise reconstruction loss assumes a precise alignment to provide an informative feedback signal. However, the online registration head will typically only succeed if a good contrast between structures is being predicted. In other words, there is a cross-dependency of accurate pixel reconstructions and producing a high contrast over training iterations, which may result in a conflict during training. This can be seen in Fig. 9, where a Gram+Reg model focussing on pixel reconstruction only made heavily blurred predictions that do not provide structural details for cell-precise registration. Adding a style loss component leads to more pronounced contours while simultaneously achieving higher pixel reconstruction accuracy (Tab. 3) through online registration. This underpins the importance of the additional style loss, which does not depend on perfect alignment and guides the training to produce structures.
@@
Apparently, the choice for Gram or GAN as style loss should be considered in the light of the application. Training with Gram loss led to more accurate cell localization and pixel values (Tab. 2). However, it was affected by staining intensity inhomogeneities, resulting in blurry cells (Fig. 8). GAN loss, on the other hand, produced higher contrast in cells but with lower accuracy. This could be a consequence of the Gram loss referring to the specific target image, while the GAN loss refers to the average scoring of the discriminator across the distribution, less than the precision with respect to an individual image.
@@
Throughout all models, we observed occasional staining inhomogeneities in the predicted virtual stainings. This effect was particularly pronounced in the form of stripe patterns in the primary motor cortex ROI (Fig. 5). Cell body membranes and blood vessel walls pose hard edges in the cortex, resulting in diffraction patterns even brighter than the background field. Varying of the focus level of the LMP-1 microscope allows optical scanning of the Poisson spots induced by diffraction. Cell bodies are highlighted with higher contrast by shifting and lowering the focal plane to the top and the bottom of the tissue by approximately $\pm$30 µm (Fig. 11). This means that the ability to predict cells from 3D-PLI measurements is sensitive to the focus level of the microscope, and it is difficult to generalize to unseen focus levels that were not part of the model training.
During the mounting process of the 60 µm thick sections in 20% glycerol solution on the object slide the gray matter is more prone to swelling than the more compact white matter, leading to tissue expansion, mechanical stress and tape flutter especially in the cortex due to its higher flexibility. The free-floating tissue between object slide and coverslip shapes waves of up to 30 µm in amplitude which affects the focus level and the light incidence on the tissue, hence the effective birefringence, modifying the polarization state of the transmitted light. Therefore, the stripe-like inhomogeneities in the virtual staining can be explained by such variations in focus levels within the same section.
An effective countermeasure against waves in the tissue is to weight the open coverslips after embedding overnight. In addition, the acquisition of multiple focus levels for each section could potentially provide missing information. For the given dataset, however, these potential solutions were not applicable, and measurements could not be easily repeated. In addition, the acquisition of multiple focus levels would significantly increase the measurement effort. Nevertheless, it will be an important consideration for future investigations to include multiple focus levels in each measurement to improve cell extraction.
While most of the larger cells (>100 µm² in-plane size) could be localized in 3D-PLI parameter maps, smaller cells were often missed or misplaced in the predictions. This is expected as smaller cell structures are often dominated by a stronger signal of intersecting nerve fibers or overlay each other due to the high section thickness of 60 µm required for 3D-PLI. The predictions also showed a stronger arrangement of cells into cortical columns than in the target images. The presence or absence of this arrangement is one of the criteria used to identify cortical areas, which might impair downstream interpretation of the predicted architecture. It may be attributed to the pronounced signal of radial fibers in 3D-PLI, which are used by the virtual staining model to infer potential cell locations. Further investigations are necessary to understand this effect.
Our ablation experiments in Sec. 3.2, Sec. 3.3 and Sec. 3.4 support the design decisions of the proposed virtual staining method. They showed an overall good reconstruction and localization of individual cell bodies. But, of course, the virtual stainings can not yet fully replace real measurements. They occasionally include staining inhomogeneities, miss smaller cells, or introduce implausible cytoarchitectonic features, such as artificial columnar arrangements of cells. In contrast, extracting GLI profiles from the predictions demonstrated that the laminar cell organization can be reconstructed quite well. However, further investigations with larger datasets spanning multiple brain areas are necessary to thoroughly assess their reliability.
@@
The proposed methodology might be further improved by extending the amount of training sections, brain regions and focus levels in 3D-PLI acquisition. We also expect that the expressiveness of the Gram matrix representation could be enhanced by replacing the VGG encoder, pre-trained on ImageNet, with a domain-specific model, which is powerful on microscopy data. This is technically motivated by the observation that higher-level features in ImageNet-pretrained networks, such as VGG, are optimized for object semantics in natural images (e.g., faces, animals, vehicles), which are irrelevant in the context of fine-grained texture characteristic of microscopy images Stuckner et al., (2022). In contrast, microscopy-specific encoders are better suited to capture such low- to mid-level textural features and have been shown to improve clustering Oberstrass et al., (2024), segmentation Stuckner et al., (2022), and the evaluation of generative models Kropp et al., (2024) over those trained on ImageNet. Therefore, computation of Gram matrices using domain-specific encoders Spitzer et al., (2017); Schiffer et al., 2021b ; Schiffer et al., 2021a or emerging foundation models for histology Tran et al., (2024) might further enhance representation of cytoarchitectural characteristics.
@@ 5 Conclusions
Motivated by previous observations that larger cells are encoded in 3D-PLI parameter maps alongside fiber orientations Zeineh et al., (2017), we introduced a deep learning model for transforming 3D-PLI maps into virtual Cresyl violet cell body stainings. This approach enables joint visualization of fiber tracts and cell bodies in the same tissue. Compared to real post-staining, the model offers a scalable alternative that avoids manual labor, limited only by its generalization to unseen architectural patterns. Validation against post-stained test sections confirms that our virtual stainings reliably localize most of the larger cell bodies (>100 µm² in-plane size).
A central contribution of our approach is the integration of an online registration head during training. This component eliminates the need for explicit multimodal registration, which is commonly required in virtual staining pipelines Rivenson et al., (2020); de Haan et al., (2021); Yang et al., (2022). It is a simple but highly efficient add-on that can be combined with various loss formulations, leveraging model-estimated landmarks, to continuously refine the alignment over time.
The method enables accurate prediction of cell bodies from 3D-PLI and successful adaptation to the appearance of real Cresyl violet stainings. As such, it expands the usability of 3D-PLI in large-scale data settings, allowing virtual staining at scale. While such synthetic data cannot and should not replace real histological measurements, it offers promising opportunities in downstream analysis. Potential applications include cross-modal image registration to align real Cresyl violet and 3D-PLI images, disentangling signal sources to identify voxels dominated by cell bodies (e.g., for improved tractography), and missing data imputation in serial section stacks. Especially in interleaved modalities, this could enable the reconstruction of complete datasets. Of course, such applications require careful quality control and distinction between synthetic and real measurements.
@@
The ROIs employed for the training and testing of the models in this study, along with a selection of model predictions, are available in our central institutional repository Oberstrass et al., (2025).
@@
Conflict of Interest Statement
@@ Appendix A Appendix
Figure 11: Virtual staining quality varies with the focus level of the LMP-1 due to differences in the captured diffraction patterns. Example images show the vervet entorhinal cortex with zoom-ins to the architecture of cortical layers II-III. 3D-PLI inputs are represented by the transmittance $I_{T}$, retardation $\sin\delta$ (scaled with gamma correction) and fiber orientation in HSV color space (hue: direction $\varphi$; saturation, brightness: retardation $\sin\delta$). The same Gram+Reg model is used to predict virtual stainings from a focus level manually adjusted to the tissue center (+0 µm) and 30 µm above and below. A lower focus level of -30 µm achieves the highest consistency in staining intensity and produces the most cells. Higher focus levels introduce an increasing amount of staining artifacts.

### v2

@@ 3.5 Influencing factors on model predictions
Figure 10: Reliability of predicted cell instances increases with their size and decreases with a stronger retardation. We report F1 scores between detected cell instances by a CPN model in the predicted and target stainings. F1 scores are computed for intervals of (A) the segmented in-plane cell body area and (B) smoothed retardation maps. Retardation values are sampled at each cell location. Each bar represents the average over four independently trained models. Error bars show standard deviation.
@@
Larger cells are expected to be more pronounced in 3D-PLI parameter maps compared to smaller cells that may be overshadowed by other tissue components such as nerve fibers. To quantify the reliability of predicted cells in the virtual staining in relation to their size, we compute F1 scores for multiple bins of 2D in-plane cell sizes. Obtaining F1 scores for each bin requires computation of true and false positives, as well as false negatives through matching of predicted and target cells with a minimum IoU threshold of 0.3. We modify the computation of true and false positives by matching only predicted cells within that range with all cells in the target image. To count false negatives, target cells within that range are matched with all cells in the prediction. Results in Fig.˜10A show that reconstruction of smaller cells < 50 µm² has much lower F1 scores, below 20.0, throughout all methods. With increasing 2D cell size, they can be identified more accurately. The Gram+Reg model shows highest F1 scores across all cell sizes and smallest variation between four independently trained models, measured as standard deviation.
Birefringent tissue, such as myelinated nerve fibers, can obscure signals from other components in 3D-PLI. To examine whether this effect impacts the prediction of cell instances in the virtual staining, we use retardation maps as a measure of birefringence strength. The maps are smoothed with a 10 pixels square median kernel to reduce local variance and obtain values representative of the surrounding tissue area. For each cell instance segmented by the CPN model in the virtual staining, we sample a retardation value at its center location. Values are grouped into intervals from 0.0 to 0.3 in steps of 0.05. Since we focus the analysis on gray matter, larger retardation values do not occur. As shown in Fig.˜10B, the ability of all models to reconstruct cell instances decreases significantly for cells dominated by stronger retardation signals. For Gram+Reg this effect is minimal, performing overall best.
Figure 11: Virtual staining quality varies with the focus level of the LMP-1 microscope, which influences the captured diffraction patterns. Example images show the vervet entorhinal cortex with zoom-ins to the architecture of cortical layers II-III. 3D-PLI inputs are represented as transmittance, retardation (scaled with gamma correction) and fiber orientation in HSV color space (hue: fiber direction; saturation/brightness: retardation). The same Gram+Reg model is used to predict virtual stainings from a focus level adjusted to the tissue center (+0 µm) and 30 µm above and below.
We observe that the virtual staining intensities are not always homogenous within generated images. Such effects can especially be observed as cloud-like patterns in Fig.˜5 for the predicted virtual staining of ROIs showing the primary motor cortex and subcortical nuclei. These variations coincide with variations in the diffraction pattern of the associated transmittance maps, which in turn can be influenced by the focus level of the LMP-1 microscope. To better quantify this relationship, we apply a Gram+Reg model to predict virtual stainings from an example image showing the vervet entorhinal cortex, captured at three focus levels: centered on the tissue (+0 µm), -30 µm and +30 µm. Results are summarized in Fig.˜11. We observe the highest consistency in staining intensity at the lowest focus level of (-30 µm). Higher focus levels introduce an increasing amount of staining artifacts.
@@ 4 Discussion
Our ablation experiments in Sec.˜3.2, Sec.˜3.3 and Sec.˜3.4 support the design decisions of the proposed virtual staining method. In Sec.˜3.5 we investigated several influencing factors on the predictions. In the following, we want to discuss the design decisions of the proposed virtual staining method and assess its reliability for downstream analysis.
@@
We introduced an online registration head capable of approximating smooth nonlinear deformations, given sufficiently small patch sizes for training. Its accuracy is inherently influenced by the choice of the registration metric. We identified MSE in the Fourier domain through multiple applications of cross-correlation as superior to other typically used metrics in Fourier-based image registration, including PC. This is a notable observation, as PC is a popular choice due to its robustness to intensity variations and frequency-dependent noise tong2019. However, during early training stages, model predictions are often blurred. The network tends to learn coarse, low-frequency structures first. Additionally, random weight initialization introduces noise in the first steps. MSE appears more resilient to such blur and frequency-independent noise, leading to more accurate predictions (Fig.˜6). This behavior can be explained by PC’s reliance on strong high-frequency spectral peaks, which are diminished by blur pedone2013. MSE, in contrast, maintains sensitivity across the frequency spectrum and can better handle smooth, low-detail inputs that dominate during initial training.
@@
The applied pixel-wise reconstruction loss assumes a precise alignment to provide an informative feedback signal. However, the online registration head will typically only succeed if a good contrast between structures is being predicted. In other words, there is a cross-dependency of accurate pixel reconstructions and producing a high contrast over training iterations, which may result in a conflict during training. This can be seen in Fig.˜9, where a Gram+Reg model focussing on pixel reconstruction only made heavily blurred predictions that do not provide structural details for cell-precise registration. Adding a style loss component leads to more pronounced contours while simultaneously achieving higher pixel reconstruction accuracy (Tab.˜4) through online registration. This underpins the importance of the additional style loss, which does not depend on perfect alignment and guides the training to produce structures. Still, some manual pre-alignment was required to keep the registration error of training data within 70 pixels (91 µm) and ensure sufficient structural overlap between training patches. This limited the amount of available training data, as larger registration errors, such as those from the initial affine alignment of whole sections, were too large for the online registration to succeed.
@@
Apparently, the choice for Gram or GAN as style loss should be considered in the light of the application. Training with Gram loss led to more accurate cell localization and pixel values (Tab.˜2). However, it was affected by staining intensity inhomogeneities, resulting in blurry cells (Fig.˜8). GAN loss, on the other hand, produced higher contrast in cells but with lower accuracy. This could be a consequence of the Gram loss referring to the specific target image, while the GAN loss refers to the average scoring of the discriminator across the distribution, less than the precision with respect to an individual image.
@@
We expect that the expressiveness of the Gram matrix representation could be enhanced by replacing the VGG encoder, pre-trained on ImageNet, with a domain-specific model, which is powerful on microscopy data. This is technically motivated by the observation that higher-level features in ImageNet-pretrained networks, such as VGG, are optimized for object semantics in natural images (e.g., faces, animals, vehicles), which are irrelevant in the context of fine-grained texture characteristic of microscopy images stuckner2022. In contrast, microscopy-specific encoders are better suited to capture such low- to mid-level textural features and have been shown to improve clustering oberstrass2024a, segmentation stuckner2022, and the evaluation of generative models kropp2024 over those trained on ImageNet. Therefore, computation of Gram matrices using domain-specific encoders spitzer2017; schiffer2021; schiffer2021a or emerging foundation models for histology tran2024 might further enhance representation of cytoarchitectural characteristics.
Throughout all models, we observed occasional staining inhomogeneities in the predicted virtual stainings. An example of local-scale variation is the inhomogeneity observed in Fig.˜8, which may be attributed to the tiling strategy used during model inference to transfer whole sections. Since the model lacks a global view of the section, it may reproduce variations in the training data at the local patch level. At a larger scale, this effect was pronounced in the form of stripe patterns, for example in the primary motor cortex ROI (Fig.˜5). These patterns typically coincide with inhomogeneities in the 3D-PLI transmittance maps. During the mounting process of the 60 µm thick sections in 20% glycerol solution on the object slide the gray matter is more prone to swelling than the more compact white matter, leading to tissue expansion, mechanical stress and tape flutter especially in the cortex due to its higher flexibility. The free-floating tissue between object slide and coverslip shapes waves of up to 30 µm in amplitude. These waves affect the focus level and the light incidence on the tissue, hence the effective birefringence, modifying the polarization state of the transmitted light. This caused the observed stripe patterns in the transmittance maps and very likely contributed to the staining inhomogeneities.
Cell body membranes and blood vessel walls pose hard edges in the cortex, resulting in diffraction patterns even brighter than the background field. Varying of the focus level of the LMP-1 microscope allows optical scanning of the Poisson spots induced by diffraction. Cell bodies are highlighted with higher contrast by shifting and lowering the focal plane to the top and the bottom of the tissue by approximately $\pm$30 µm (Fig.˜11). The ability to predict cells from 3D-PLI measurements was shown to be sensitive to variations in the focus level. An effective countermeasure against waves in the tissue, to maintain a constant level across the entire section, is to weight the open coverslips after embedding overnight. In addition, the acquisition of multiple levels for each section could potentially provide missing information. For the given dataset, however, these potential solutions were not applicable, and measurements could not be easily repeated. Since shifting the focus toward the tissue surface degrades optical resolution, a center focus level is used by default in 3D-PLI for an optimal representation of nerve fibers. The acquisition of multiple focus levels of $\pm$30 µm to increase the visibility of cell bodies would significantly increase the measurement effort. Nevertheless, it will be an important consideration for future investigations to include multiple focus levels in each measurement to improve cell extraction.
It must be noted that shrinking and swelling of the tissue also affects the measured areas of individual 2D segmented cell bodies. Tissue deformation in histology is complex, involving anisotropic effects and differences between cellular and extracellular structures dorph-petersen2001; west2013, which cannot be fully captured by the global shrinkage factor estimated in Sec.˜2.1.3. At the same time, there is only sparse data available that quantifies such effects, and the precise conditions of histological processing differ between each other, and it can be assumed that such differences influence the specific shrinkage values. In paraffin-embedded human sections, a mean difference of 9% between gray and white matter shrinkage has been reported kretschmann1982. As frozen sections generally show less in-plane deformation than paraffin sections (west2013), we treat this value as an upper bound for our acqusition. Since gray matter contains relatively more cell bodies than white matter, similar shrinkage of gray and white matter suggests that shrinkage of cell bodies and neuropil should also not differ significantly amunts1999. By performing non-linear co-registration of Cresyl violet to 3D-PLI and applying a subsequent global correction of cell sizes to account for swelling effects from MRI to 3D-PLI, we assume that major shrinkage effects are compensated. However, residual local variations or relative slice-to-slice differences may persist and will require a more systematic approach in future work.
While most of the larger cells (>100 µm² in-plane size) could be localized in 3D-PLI parameter maps, smaller cells were often missed or misplaced in the predictions. This is expected as smaller cell structures are often dominated by a stronger signal of intersecting nerve fibers or overlay each other due to the high section thickness of 60 µm required for 3D-PLI. In our experiments, we observed that stronger birefringence, corresponding to higher retardation values, resulted in reduced capability of models to reconstruct cell instances. Since birefringence increases with myelin density and homogeneity, predictions in regions with densely packed myelinated nerve fibers, such as white matter, are thus less reliably captured by the virtual stainings. In white matter, detecting cell bodies may even be infeasible, as the predominant oligodendrocytes form the myelin sheath and are therefore indistinguishable from it. In gray matter, this limitation may explain the stronger arrangement of cells into cortical columns in the predictions compared to the target images in Fig.˜8, as the strong signal of myelinated radial fibers can obscure underlying cell instances. The presence or absence of this arrangement is one of the criteria used to identify cortical areas, which might impair downstream interpretation of the predicted architecture.
As shown by the differences in F1 values in Tab.˜5, cells were reconstructed more accurately in isocortical areas (motor cortex and temporal cortex) than in the hippocampus or the subcortical nuclei (putamen and globus pallidus). The lower F1 in the putamen compared to the motor and temporal areas can be attributed to its neuronal composition and the fundamentally different spatial arrangement of cell bodies in cortical versus subcortical regions, as well as their relation to fiber bundles. The putamen is composed of medium-sized spiny neurons and relatively small interneurons arranged around and between bundles of myelinated fibers, which course through the putamen and give rise to its characteristic striated appearance heilbronner2025. In this region, two challenges reduce predictability: small interneurons are difficult to detect with our approach, and fiber bundles may intersect or obscure cell bodies, making them harder to reconstruct. Moreover, the cell bodies of spiny neurons can resemble fiber bundles oriented perpendicular to the section plane, appearing as dark patches in the transmittance image. The globus pallidus contains relatively large but sparsely distributed neurons, which are embedded in a dense matrix of multidirectionally oriented myelinated fibers heilbronner2025. Here, predictability is reduced by the number and complex spatial organization of intersecting myelinated fiber bundles and blood vessels. The low F1 value in patches extracted from the CA1-CA3 regions of the hippocampus is more difficult to explain, though it may result from the fact that they cover both the pyramidal layer, with its numerous and relatively large pyramidal cells, and the radiatum layer, which contains sparsely packed and relatively small interneurons zhao2025, which are difficult to detect with our method.
Our Gram+Reg model produced larger cells at expected positions and generated images with a plausible appearance, making virtual stainings useful for cross-modality registration - at present the most relevant application. They also allowed the application of cytoarchitectonic tools such as cell segmentation or computation of GLI profiles for cortical layer characterization on 3D-PLI images. As a future perspective, the ability to detect cell bodies in 3D-PLI may enhance the computation of 3D fiber orientations within gray matter, which is a prerequisite for fiber tractography. This could be achieved by improving the estimation of myelination, for example through the identification of voxels dominated by cell bodies, and potentially also support the localization of axon terminals. However, virtual stainings occasionally contained artifacts, including staining inhomogeneities, omission of smaller cells, or the introduction of implausible cytoarchitectonic features. These issues may reflect biases in the training data, model-related artifacts, or missing cellular signatures in the 3D-PLI parameter maps. The virtual staining allowed the identification by a neuroanatomist (N.P.-G.) of the borders between the CA1, CA2, and CA3 regions of the hippocampus, the border between the primary motor and primary somatosensory areas, or the border between the retrosplenial cortex and cingulate area 23. However, it was not possible to identify the border between the core and lateral belt auditory areas, because it is characterized by differences in the packing density of small pyramids hackett2001, which currently cannot be reliably predicted. Therefore, we do not yet consider the method sufficiently robust for reliable cytoarchitectonic brain area mapping. Nevertheless, we are convinced that expanding the number of training sections, brain regions, and focus levels across brains and species can improve model performance and generalizability. While such expansion would require re-training, each additional sample would teach the model novel architectural patterns, improving its robustness across domains.
@@ 5 Conclusions
Motivated by previous observations that larger cells are encoded in 3D-PLI parameter maps alongside fiber orientations zeineh2017, we introduced a deep learning model for transforming 3D-PLI maps into virtual Cresyl violet cell body stainings. This approach enables joint visualization of fiber tracts and cell bodies in the same tissue. Compared to real post-staining, the model may offer a scalable alternative that avoids manual labor.
A central contribution of our approach is the integration of an online registration head during training. This component eliminates the need for explicit, pixel-accurate multimodal registration, which is commonly required in virtual staining pipelines rivenson2020; dehaan2021; yang2022. It is a simple but highly efficient add-on that can be combined with various loss formulations, leveraging model-estimated landmarks, to continuously refine the alignment over time.
The developed method enables localization of most of the larger cell bodies in gray matter (>100 µm² in-plane size) from 3D-PLI and a successful adaptation to the appearance of real Cresyl violet stainings. As such, it expands the usability of 3D-PLI in large-scale data settings, allowing virtual staining at scale. While such synthetic data cannot and should not replace real histological measurements, it offers promising opportunities in downstream analysis. Although we see the present method not yet robust enough for cytoarchitectonic mapping, potential applications include cross-modal image registration to align real Cresyl violet and 3D-PLI, performing cell segmentation in 3D-PLI images, or missing data imputation in serial section stacks. Especially in interleaved modalities, this could enable the reconstruction of complete datasets. Of course, such applications require careful quality control, a clear separation and demarcation of synthetic data from real measurements, and careful interpretation of derived results.
@@
ROIs employed for the training and testing of the models in this study, along with a selection of model predictions, are available in our central institutional repository oberstrass2025dataset. In addition, the repository includes whole-slide predictions for test section 559 along with corresponding 3D-PLI modalities.
@@
Declaration of Competing Interests
@@ Appendix A Appendix
Table 5: Specific quantitative results per test ROI, corresponding to averages reported in Tab.˜2. The deviation is reported as standard error over four independent trainings with different random seeds. Arrows indicate the direction of better performance ($\uparrow$ higher is better, $\downarrow$ lower is better). Best scores per ROI in bold.
ROI Method MI $\uparrow$ RMSE $\downarrow$ SSIM $\uparrow$ F1 $\uparrow$
Motor cortex Gram 0.061 $\pm$ 0.013 33.9 $\pm$ 1.0 0.306 $\pm$ 0.024 24.8 $\pm$ 3.5
Gram+Reg 0.168 $\pm$ 0.003 29.2 $\pm$ 0.6 0.446 $\pm$ 0.002 42.9 $\pm$ 0.3
GAN 0.117 $\pm$ 0.016 34.7 $\pm$ 1.1 0.268 $\pm$ 0.010 16.8 $\pm$ 1.0
GAN+Reg 0.165 $\pm$ 0.015 29.6 $\pm$ 1.0 0.394 $\pm$ 0.013 36.3 $\pm$ 2.1
Hippocampus Gram 0.184 $\pm$ 0.015 34.5 $\pm$ 0.9 0.340 $\pm$ 0.013 19.4 $\pm$ 2.3
Gram+Reg 0.303 $\pm$ 0.003 31.1 $\pm$ 0.5 0.448 $\pm$ 0.001 31.4 $\pm$ 0.4
GAN 0.078 $\pm$ 0.016 44.3 $\pm$ 3.1 0.262 $\pm$ 0.016 6.9 $\pm$ 1.3
GAN+Reg 0.151 $\pm$ 0.036 32.1 $\pm$ 0.6 0.373 $\pm$ 0.027 24.9 $\pm$ 1.9
Temporal cortex Gram 0.077 $\pm$ 0.018 33.9 $\pm$ 0.8 0.319 $\pm$ 0.040 31.7 $\pm$ 4.5
Gram+Reg 0.236 $\pm$ 0.003 29.1 $\pm$ 0.4 0.497 $\pm$ 0.004 52.7 $\pm$ 0.4
GAN 0.051 $\pm$ 0.006 43.1 $\pm$ 5.4 0.238 $\pm$ 0.019 18.3 $\pm$ 2.7
GAN+Reg 0.192 $\pm$ 0.015 30.1 $\pm$ 1.5 0.460 $\pm$ 0.017 49.6 $\pm$ 1.3
Subcortical nuclei Gram 0.122 $\pm$ 0.007 32.0 $\pm$ 0.6 0.290 $\pm$ 0.016 15.7 $\pm$ 2.2
Gram+Reg 0.190 $\pm$ 0.004 29.2 $\pm$ 0.4 0.388 $\pm$ 0.001 29.4 $\pm$ 0.5
GAN 0.138 $\pm$ 0.028 33.0 $\pm$ 0.8 0.245 $\pm$ 0.013 11.4 $\pm$ 0.9
GAN+Reg 0.132 $\pm$ 0.014 29.9 $\pm$ 0.7 0.330 $\pm$ 0.014 20.9 $\pm$ 1.6